CrCl3 + AgNO3 > AgCl + Cr(NO3)3
A 0.720g sample of impure CrCl3 was dissolved in water and required 43.5 mL of 0.115 M AgNO3 solution to react with the CrCl3 in the sample. What is the mass percent of CrCl3 in the sample?
1
Expert's answer
2019-10-08T07:43:20-0400

The reaction occurs by the following equation:

"CrCl_3 + 3AgNO_3 \rightarrow 3AgCl + Cr(NO_3)_3"

The amount of silver nitrate required for the reaction can be found:

"n(AgNO_3) = c(AgNO_3) \cdot V(AgNO_3) = 0.115 \frac{mol}{L} \cdot 0.0435 L = 0.005 mol"

According to stoichiometry of the reaction the amount of CrCl3 is one third of silver nitrate:

"n(CrCl_3) = \frac{1}{3} \cdot n(AgNO_3)"

The mass of chromium chloride reacted with silver nitrate is:

"m(CrCl_3) = n(CrCl_3) \cdot M(CrCl_3) = \frac{1}{3} \cdot 0.005 mol \cdot 158.5 \frac{g}{mol} = 0.264 g"

The mass percent of chromium chloride in the sample is:

"\omega (CrCl_3) = \frac{m(CrCl_3)}{m(sample)} \cdot 100 \% = \frac{0.264 g}{0.720 g} \cdot 100 \% =36.7 \%"